OpenStudy (anonymous):

How do I solve these? log5x + log4 = 3

JoelTheBoss (joel_the_boss):

Have you ever attempted to solve logs before?

OpenStudy (alekos):

do you know the log rules?

OpenStudy (amorfide):

\[\log(a)+\log(b)=\log(ab)\] you have log(5x) you can seperate this

OpenStudy (anonymous):

We just started learning them this week and at first we were just doing the basic log rearranging of them, but I'm so lost with this. I think I missed something.

OpenStudy (amorfide):

\[\log(a)-\log(b)=\log(a/b)\]

OpenStudy (anonymous):

I don't understand. Do I plug the numbers into that? or?

OpenStudy (amorfide):

okay so if I have \[\log(4)+\log(x)=\log(4x)\] then it is okay to say I can split up log(4x) \[\log(4x)=\log(4)+\log(x)\] so if you seperate log(5x) then solve your given expression to get it in the form of log(x)=.... try this then I will help you from here

OpenStudy (anonymous):

log4 = 5x + 3?

OpenStudy (amorfide):

no

OpenStudy (amorfide):

log(5x)=log(5)+log(x) so replace it

OpenStudy (amorfide):

log(5)+log(x)+log(4)=3

OpenStudy (amorfide):

now use your rule for log(a)+log(b)=log(ab)

OpenStudy (anonymous):

Oh okay sorry

OpenStudy (amorfide):

\[\log(5)+\log(4)=\log(5 \times 4)\]

OpenStudy (amorfide):

replace it

OpenStudy (amorfide):

log(x)+log(20)=3

OpenStudy (amorfide):

now re arrange it to get log(x)=... can you do this?

OpenStudy (anonymous):

log(20x)=3?

OpenStudy (amorfide):

okay if you wanna do it that way sure

OpenStudy (amorfide):

now I need to know, if we are in base 10 or natural log

OpenStudy (anonymous):

In base 10

OpenStudy (amorfide):

now I need to know, if we are in base 10 or natural log

OpenStudy (anonymous):

Do you understand the rules of exponents?

OpenStudy (amorfide):

okay i am getting some serious issues with this website one sec

OpenStudy (anonymous):

Wio no I don't understand them that well

OpenStudy (amorfide):

\[\log_{a}(b)=Y\] \[a^{Y}=b\]

OpenStudy (anonymous):

You can leverage your own understanding of exponents to do logarithms. \[ \log5x + \log4 = 3\\ 10^{\log 5x+\log 4}=10^3\\ 10^{\log 5x}\cdot10^{\log 4} = 1000\\ (5x)(4) = 1000 \]

OpenStudy (amorfide):

@wio can you finish this? I am having some issues

OpenStudy (anonymous):

However, in this case, just using logarithm properties are just good, since you're not familiar with either.

OpenStudy (amorfide):

log(20x)=3 since we are already this far you know the base is 10 and since I posted the rule for the log where you do the base raised to the power of the answer so in this case 10^3=20x then solve for x

OpenStudy (anonymous):

The exponent property: \[ x^a\cdot x^b=x^{a+b} \]Corresponds to the logarithm property: \[ \log(a)+\log(b) = \log(a\cdot b) \]

OpenStudy (anonymous):

x=50?

OpenStudy (alekos):

that's it

OpenStudy (anonymous):

yes, also numerical solution via tool is 50 :) http://www.equationcalculator.org/?input=log+%285x%29+%2B+log+%284%29+%3D+3&submit=Calculate

OpenStudy (anonymous):

Thank you guys :)